Suppose a baseball player "pops up", that is, hits the baseball straight up into the air. The height, h, of the baseball t seconds after it leaves the bat can be calculated using a formula from physics that takes into account the initial speed of the ball and the height at which it was hi. The formula for the height of the ball in feet in h= -16t^2+64t+4
Graph the height of the baseball versus time. connect the dots with a smooth curve.
What are the coordinates of tye highest point on the graph? When does the baseball reach its maximum height? What is that height? By Answer by richwmiller(17219) 
 h= -16t^2+64t+4
using the derivative for the vertex
-32t+64=0
h= -16*2^2+64*2+4
h=-16*4+64*2+4
h=-64+128+4
h=68
t=2 h=68
or using -b/2a for the vertex
t=-64/-32=2
h=68 from above
(2,68)
At 2 seconds it reaches 68 feet.